# Grounded-VideoLLM: Sharpening Fine-grained Temporal Grounding in Video Large Language Models

### v1

@@ 6.2 Main Results
Table 4: Results on zero-shot Open-Ended VideoQA and VCG-Bench. VCG-Bench contains five aspects: Correctness of Information (CI), Detail Orientation (DO), Contextual Understanding (CU), Temporal Understanding (TU), and Consistency (CO).
@@
Grounded-VideoLLM 59.4 76.0 75.5 77.0 48.0 67.5 85.5 77.0 34.5 39.5 59.5 86.5 44.5 60.5 79.0 51.5 49.0 46.0 35.0 42.5 54.0
Table 5: Results on MVBench multi-choice question answering.
NExT-GQA (Xiao et al., 2024) is quite challenging since it requires the model to not only correctly answer questions but also provide timestamps that support the answers, highlighting the temporal reasoning capability. According to Table 3, Grounded-VideoLLM achieves the highest Acc@GQA (26.7, +2.4) and delivers comparable IoU and IoP scores to models such as SeViLA (Yu et al., 2023) and LLoVi (Zhang et al., 2023a), which use specialized grounding modules or rely on proprietary large language models (Achiam et al., 2023). The highest Acc@GQA score further demonstrates Grounded-VideoLLM’s capability in both fine-grained temporal grounding and high-level reasoning.
General Video-LLM Benchmarks. While Grounded-VideoLLM excels in fine-grained temporal grounding, we aim to ensure that it maintains robust performance in general video understanding. Therefore, we conducted a comprehensive evaluation using VCG-Bench (Maaz et al., 2024b) and MVBench (Li et al., 2024). As shown in Table 4, Grounded-VideoLLM achieves promising results in VCG-Bench, with an average score of 3.24, outperforming other Video-LLMs with temporal grounding capabilities (e.g., LITA, VTimeLLM). Notably, Grounded-VideoLLM surpasses all other Video-LLMs on the TU (Temporal Understanding) task, with a score of 3.12 (+7%), demonstrating its superior temporal understanding, which can be attributed to the two-stream architecture that can capture motion dynamics. For MVBench which provides 4,000 QA pairs spanning a wide range of scenes categorized into 20 fine-grained tasks, the results, presented in Table 5, show that Grounded-VideoLLM achieves an average score of 59.4, surpassing other Video-LLMs. Notably, it achieves top performance in several critical tasks requiring perceiving and reasoning over specific video moments, including Action Sequence (AS), Action Prediction (AP), Action Localization (AL), and State Change (SC), with scores of 76.0 (+10%), 75.0 (+26%), 59.5 (+43%), and 51.5 (+14%), respectively, demonstrating significant advancements on fine-grained video understanding.
6.3 In-Depth Analysis
Figure 4: We visualize the attention weights of the LLM when generating the temporal tokens.
Two-stream Encoding. We conduct ablations to our two-stream encoding. Specifically, we set two variants by removing the temporal stream while only retaining the spatial stream, where all frame embeddings are concatenated as the video representation: (1) w/o temporal-stream (dense) feeds $T=96$ frames with a pooling size $\sigma_{S}=4$ (36 tokens per frame), resulting in a total of $36\times 96=3456$ tokens. (2) w/o temporal-stream (sparse) feeds $T=24$ frames with a pooling size $\sigma_{S}=2$ (144 tokens per frame), also resulting in a total of $144\times 24=3456$ tokens. Both variants have a close number of tokens for video representation compared to our two-stream encoding ($326412\times(144+\frac{96}{12}\times 16)=$ 3264 tokens for $\textbf{F}_{Vid}\in\mathbb{R}^{K\cdot(N_{S}+\frac{T}{K}\cdot N_{T})\times D}$). Table 6 shows that both variants result in a significant performance drop on temporal grounding tasks. Interestingly, the dense frame variant performs slightly better than the sparse frame variant, suggesting that dense temporal information is more critical for grounding tasks than spatial details. Our two-stream architecture strikes a balance by effectively capturing dense motion dynamics while maintaining essential appearance details.
Model C-STA ANet-G ANet-Cap
mIoU mIoU SODA_c METEOR
Grounded-VideoLLM 36.8 36.1 6.0 6.8
w/o temporal-stream (sparse) 30.4 ($\downarrow$ 6.4) 28.0 ($\downarrow$ 8.1) 4.9 ($\downarrow$ 1.1) 5.5 ($\downarrow$ 1.3)
w/o temporal-stream (dense) 34.3 ($\downarrow$ 2.5) 29.2 ($\downarrow$ 6.9) 5.4 ($\downarrow$ 0.6) 6.2 ($\downarrow$ 0.6)
w/o temporal token alignment 27.5 ($\downarrow$ 9.3) 23.1 ($\downarrow$ 13.0) 4.7 ($\downarrow$ 1.3) 6.4 ($\downarrow$ 0.4)
Table 6: Impact of two-stream encoding and alignment stage.
Temporal Token Alignment. We further investigate the role of temporal tokens by ablation of the second training stage, Temporal Token Alignment. Quantitative results in Table 6 reveal a sharp performance drop across all tasks, particularly in temporal sentence grounding tasks. Upon analyzing the outputs, we found that the model often produces time intervals spanning nearly the entire video (e.g., from <0> to <300>), neglecting the association between specific moments and temporal tokens, which leads to misalignment. Qualitatively, we visualize the attention weights of the last layer in the LLM to demonstrate how temporal tokens attend to corresponding video moments. Details of the visualizations are provided in Appendix A.4. As shown in Figure 4 (a), when generating the temporal token, e.g. <241> or <271>, the attention weights are higher and more focused on their corresponding video moments. Conversely, in Figure 4 (b), when the model is trained without the Temporal Token Alignment stage, the attention weights of temporal tokens become significantly dispersed across irrelevant moments. This illustrates that our multi-stage training strategy is essential for achieving proper alignment between temporal tokens and the video timeline. We also visualize the embedding distribution of temporal tokens in Appendix A.5.
Model NExT-GQA
Acc@GQA mIoP mIoU
Grounded-VideoLLM 26.7 34.5 21.1
Table 7: Impact of grounded VideoQA dataset.
Grounded VideoQA Dataset. We validate the role of our constructed grounded-VideoQA dataset by removing it from the training sets of stage-3. Since the model without training on our dataset usually generates free-form texts when asked to output the timestamps supporting the answer, we reformulate it as a temporal sentence grounding task, where we combine the predicted answer and question into a single sentence and ask the model to localize its timestamps. Table 7 suggests that there is a significant performance decrease with regard to Acc@GQA ($\downarrow$ 8.6), mIoP ($\downarrow$ 12.3), and mIoU ($\downarrow$ 8.2), from which we can conclude that our Grounded VideoQA dataset is essential to further enhance the model’s temporal reasoning capability.
7 Conclusion
We present Grounded-VideoLLM, a Video-LLM capable of fine-grained perception and reasoning over specific video moments. This is achieved through a novel model architecture that incorporates two-stream encoding for effective temporal modeling, along with the temporal tokens for efficient timestamp representation. We employ a multi-stage training scheme, starting with an image-based MLLM and progressively equipping it with fine-grained temporal grounding capabilities. Additionally, we curate a grounded-VideoQA dataset to further enhance the model’s temporal reasoning ability. Extensive experiments demonstrate that Grounded-VideoLLM not only excels in video temporal grounding tasks but also performs strongly on general video understanding benchmarks.
Appendix A Appendix
@@ A.1 More Implementation Details
Phi3.5-Vison-Instruct (Microsoft, 2024) consists of a CLIP style ViT image encoder (Radford et al., 2021), an MLP projector $f(\cdot)$, and the large language model Phi3.5-mini-3.8B (Abdin et al., 2024). Each video is sampled as a sequence of $T=96$ frames, which are evenly divided into $K=12$ segments. For the spatial stream encoded by the ViT in Phi3.5-V, we adopt a higher resolution 336$\times$336, but a lower resolution 224$\times$224 for the temporal stream encoded by InternVideo-2.
We set the pooling size $\sigma_{S}$ to be 2 while $\sigma_{T}$ to be 4 respectively. For the spatial stream, each frame takes up $24\times 24=576$ tokens before while $12\times 12=144$ tokens after pooling. For the temporal stream, each frame takes up $16\times 16=256$ tokens before while $4\times 4=16$ tokens after pooling. Therefore, we have an overall of $K\times(144+\frac{T}{K}\times 16)=3264$ tokens in total.
During training, we use the AdamW optimizer with a cosine learning rate decay and set the learning rate as 2e-5 and 1e-3 for projector $f(\cdot)$ and $g(\cdot)$ in stage-1. During stage-2 and stage-3, we set the learning rate for both projectors and word embeddings as 2e-5, while 2e-4 for LoRA parameters ($r$ = 128 and $\alpha$ = 256). All experiments are conducted on 8 NVIDIA A100/H800 GPUs.
@@ A.2 Instructions for Each Task
A.3 Evaluation Process
For the evaluation of the temporal sentence grounding task, we directly input the prompt [”At which time interval in the video can we see $<query>$ occurring?”] in Table 9 to get the response ["From $<start>$ to $<end>$"], and then calculate the predicted timestamps with the Equation (3) to get the IoU metrics.
For the evaluation of the dense video captioning task, we directly input the prompt [”Detect and report the start and end timestamps of activity events in the video, along with descriptions.”] in Table 9 to get the response ["From $<start_{1}>$ to $<end_{1}>$, $<caption_{1}>$. From $<start_{2}>$ to $<end_{2}>$, $<caption_{2}>$. $\cdots$"], and then calculate the SODA_c (Fujita et al., 2020) and Meteor scores (Banerjee & Lavie, 2005).
For the evaluation of the visually-grounded VideoQA task, we adopt a two-round conversation evaluation as follows:
@@ A.3 Evaluation Process
USER: Question: $<question>$. Options: $<options>$.
ASSISTANT: Answer: $<answer>$. Round-2:
@@
ASSISTANT: From $<start>$ to $<end>$.
@@ A.4 Visualization Process
We visualize the attention weights from the last layer of the LLM during the generation of a new temporal token. Since the full video representation consists of a total of $K\times(N_{S}+\frac{T}{K}\times N_{T})$ tokens—where $T$, $K$, $N_{S}$, and $N_{T}$ denote the number of frames, number of segments, number of tokens per frame for the spatial stream, and number of tokens per frame for the temporal stream, respectively—we obtain an attention weight vector with the shape $[K\times(N_{S}+\frac{T}{K}\times N_{T}),1]$. First, we discard the spatial stream portion, focusing only on the temporal information, which results in a new vector with the shape $[K\times\frac{T}{K}\times N_{T},1]$. We then reshape this vector to the form $[T,N_{T},1]$ and average it along the spatial dimension, yielding $[T,1]$, which represents the final attention weights corresponding to each frame when generating a new token.
A.5 Distribution of Temporal Tokens
@@
We visualize the embeddings of the $M=300$ temporal tokens to investigate their distribution in embedding space. We employ PCA (Abdi & Williams, 2010) to reduce the dimensionality of the temporal tokens to 1D, 2D, and 3D representations. For all reductions, we use the reduced values as coordinates, incorporating a gradient color scheme in which the color of the data points changes progressively with the token index, as illustrated in Figure 5. Our observations reveal that temporal tokens with similar indices tend to cluster together, exhibiting a continuous transition from tokens with smaller indices (light colors) to those with larger indices (darker colors).
Figure 5: Visualization of temporal tokens with PCA.
@@
System: You are a good question generator. I need your help in generating question-answer pairs pertaining to the visual event descriptions. I have a video and I will provide you with descriptions of certain segments and their corresponding timestamps within this video. You need to consider these segments comprehensively based on the given description and timestamps, and select one segment which you think can provide a HIGH-QUALITY QUESTION. Based on the description of that segment, ask a question related to that segment, as well as one correct answer. Both the proposed answer and question should be consistent with the content of the give description. BE CAREFUL! Your proposed questions and answers should follow these rules: (0) Avoid choosing the segment spanning across the whole video. (1) The question you raised should include causal and temporal relationships as much as possible. Question types should be diverse including WHY, HOW, WHAT, WHERE, etc. (2) NEVER involve anything that is not covered in the given descriptions. (3) The answer should NEVER appear in your question. (4) Your answer should be a phrase no more than 7 words. Keep your answers concise and accurate. Demonstrations: User: video duration: 82.73 seconds segment-1: [0.83, 19.86] A young woman is seen standing in a room and leads into her dancing. segment-2: [17.37, 60.81] The girl dances around the room while the camera captures her movements. segment-3: [56.26, 79.42] She continues dancing around the room and ends by laying on the floor. Response: chosen segment: segment-3 segment timestamps: [56.26, 79.42] question: What did the girl do after she ended dancing? answer: lay on the floor $\cdots$ (other in context demonstrations) $\cdots$
@@
Prompts for Video-Caption Alignment: (1) ”Describe the following video concisely.”, (2) ”Provide a brief description of the given video clip.”, (3) ”Offer a succinct explanation of the footage presented.”, (4) ”Summarize the visual content of the following video.”, (5) ”Give a short and clear explanation of the subsequent video clip.”, (6) ”Share a concise interpretation of the video provided.”, (7) ”Present a compact description of the clip’s key features.”, (8) ”Relay a brief, clear account of the video shown.”, (9) ”Render a clear and concise summary of the video below.”, (10) ”Write a terse but informative summary of the following video clip.” Prompts for Temporal Sentence Grounding: (1) ”When does $<query>$ happen in the video?”, (2) ”At what time does the occurrence $<query>$ take place in the video?”, (3) ”During which part of the video does $<query>$ occur?”, (4) ”When in the video does the $<query>$ incident occur?”, (5) ”At which moment does $<query>$ take place in the video?”, (6) ”During which phase of the video does $<query>$ happen?”, (7) ”When does the $<query>$ event occur in the video?”, (8) ”At what time does $<query>$ occur in the video sequence?”, (9) ”When does the $<query>$ situation take place in the video?”, (10) ”At which time interval in the video can we see $<query>$ occurring?” Prompts for Dense Video Captioning: (1) ”Localize a series of activity events in the video, output the start and end timestamp for each event, and describe each event with sentences.”, (2) ”Detect and report the start and end timestamps of activity events in the video, along with descriptions.”, (3) ”Pinpoint the time intervals of activity events in the video, and provide descriptions for each event.”, (4) ”Can you compile a list of the activities and their timestamps featured in the video?”, (5) ”I need you to scrutinize the video and catalog every event it contains, along with the timestamps.” Prompts for Temporal Referring: (1) ”What is happening from $<start>$ to $<end>$?”, (2) ”What is taking place between $<start>$ and $<end>$?”, (3) ”What events unfold between $<start>$ and $<end>$?”, (4) ”What is happening during the period from $<start>$ to $<end>$?”, (5) ”What occurs between $<start>$ and $<end>$?”, (6) ”What is going on from $<start>$ to $<end>$?”, (7) ”How do things progress from $<start>$ to $<end>$?”, (8) ”Can you describe what happens from $<start>$ to $<end>$?”, (9) ”Describe the events occurring between $<start>$ and $<end>$.”, (10) ”Narrate the actions that unfold from $<start>$ to $<end>$.”

### v2

@@ Appendix A Appendix
Table 8: Results on VCG-Bench. VCG-Bench contains five aspects: Correctness of Information (CI), Detail Orientation (DO), Contextual Understanding (CU), Temporal Understanding (TU), and Consistency (CO).
@@
Grounded-VideoLLM 60.0 75.0 75.5 83.0 50.0 63.0 88.0 77.0 37.0 41.5 50.0 91.5 45.0 57.5 82.0 49.5 55.0 45.5 32.0 42.0 59.0
Table 9: Results on MVBench. MVBench contains 20 aspects: Action Sequence (AS), Action Prediction (AP), Action Antonym (AA), Fine-grained Action (FA), Unexpected Action (UA), Object Existence (OE), Object Interaction (OI), Object Shuffle (OS), Moving Direction (MD), Action Localization (AL), Scene Transition (ST), Action Count (AC), Moving Count (MC), Moving Attribute (MA), State Change (SC), Fine-grained Pose (FP), Character Order (CO), Egocentric Navigation (EN), Episodic Reasoning (ER), Counterfactual Inference (CI), and the average of all 20 metrics (Avg).
Scale R@0.3 R@0.5 R@0.7 mIoU R@0.3 R@0.5 R@0.7 mIoU SODA_c METEOR
Grounded-VideoLLM 4B 70.2 55.9 33.2 49.4 64.9 47.8 30.4 47.2 6.6 6.5
Table 10: More results on temporal sentence grounding and dense video captioning tasks. We incorporate a subset of Charades-STA and ActivityNet-Captions into the 3rd training stage and achieve better performance.
Model Acc@GQA mIoP IoP@0.3 IoP@0.5 mIoU IoU@0.3 IoU@0.5
Grounded-VideoLLM 29.4 37.4 48.0 36.5 27.0 41.0 25.8
Table 11: Results on NExT-GQA. We incorporate a subset of Charades-STA and ActivityNet-Captions into the 3rd training stage and achieve better performance.
Figure 5: Visualization of temporal tokens with PCA.
@@ A.1 More Implementation Details
Phi3.5-Vison-Instruct (Microsoft, 2024) consists of a CLIP style ViT image encoder (Radford et al., 2021), an MLP projector $f(\cdot)$, and the large language model Phi3.5-mini-3.8B (Abdin et al., 2024). Each video is sampled as a sequence of $T=96$ frames, which are evenly divided into $K=12$ segments. For the spatial stream encoded by the ViT in Phi3.5-V, we adopt a higher resolution $\times336$, but a lower resolution $\times224$ for the temporal stream encoded by InternVideo-2. We set the pooling size $\sigma_{S}$ to be 2 while $\sigma_{T}$ to be 4 respectively. For the spatial stream, each frame takes up $24\times 24=576$ tokens before while $12\times 12=144$ tokens after pooling. For the temporal stream, each frame takes up $16\times 16=256$ tokens before while $4\times 4=16$ tokens after pooling. Therefore, we have an overall of $K\times(144+\frac{T}{K}\times 16)=3264$ tokens in total. During training, we use the AdamW optimizer with a cosine learning rate decay and set the learning rate as 2e-5 and 1e-3 for projector $f(\cdot)$ and $g(\cdot)$ in stage-1. During stage-2 and stage-3, we set the learning rate for both projectors and word embeddings as 2e-5, while 2e-4 for LoRA parameters ($r$ = 128 and $\alpha$ = 256). All experiments are conducted on NVIDIA A100/H800 GPUs.
@@ A.2 Instructions for Each Task
A.3 Grouned-VideoQA Dataset Generation
A.4 Evaluation Process
For temporal sentence grounding, we report the metric of Intersection over Union (IoU) (Gao et al., 2017) between the timestamps predicted by the model and the ground truth, including Recall at IoU thresholds of {0.3, 0.5, 0.7} and their mean IoU. For dense video captioning, we use metrics including SODA_c (Fujita et al., 2020) which is specifically tailored for the video’s storyline, and METEOR score (Banerjee and Lavie, 2005), which is the average of traditional METEOR scores that are calculated based on matched pairs between generated events and the ground truth across IoU thresholds of {0.3, 0.5, 0.7, 0.9}. For Visually-grounded VideoQA, we calculate both the Intersection of Prediction (IoP) (Xiao et al., 2024) and Intersection of Union (IoU), and use Acc@GQA (Xiao et al., 2024) to measure the percentage of questions that are both correctly answered and visually grounded with IoP $\geq$ 0.5. The responses of Open-Ended VideoQA and VCG-Bench are evaluated by GPT-3.5 with the prompts introduced by Video-ChatGPT (Maaz et al., 2024b).
For the evaluation of the temporal sentence grounding task, we directly input the prompt [”At which time interval in the video can we see <query><query> occurring?”] in Table 13 to get the response ["From <start><start> to <end><end>"], and then calculate the predicted timestamps with the Equation (3) to get the IoU metrics.
For the evaluation of the dense video captioning task, we directly input the prompt [”Detect and report the start and end timestamps of activity events in the video, along with descriptions.”] in Table 13 to get the response ["From $start_{1}>$ to $end_{1}>$, $caption_{1}>$. From $start_{2}>$ to $end_{2}>$, $caption_{2}>$. $\cdots$"], and then calculate the SODA_c (Fujita et al., 2020) and Meteor scores (Banerjee and Lavie, 2005).
For the evaluation of the visually-grounded VideoQA task, we adopt a two-round conversation evaluation:
@@ A.4 Evaluation Process
USER: <question><question>. <options><options>.
ASSISTANT: Answer: <answer><answer>. Round-2:
@@
ASSISTANT: From <start><start> to <end><end>.
@@ A.5 Visualization Process
We visualize the attention weights from the last layer of the LLM during the generation of a new temporal token. Since the full video representation consists of a total of $K\times(N_{S}+\frac{T}{K}\times N_{T})$ tokens—where $T$, $K$, $N_{S}$, and $N_{T}$ denote the number of frames, number of segments, number of tokens per frame for the spatial stream, and number of tokens per frame for the temporal stream, respectively—we obtain an attention weight vector with the shape $K\times(N_{S}+\frac{T}{K}\times N_{T}),1]$. First, we discard the spatial stream portion, focusing only on the temporal information, which results in a new vector with the shape $K\times\frac{T}{K}\times N_{T},1]$. We then reshape this vector to the form $T,N_{T},1]$ and average it along the spatial dimension, yielding $[T,1]$, which represents the final attention weights corresponding to each frame when generating a new token.
A.6 Distribution of Temporal Tokens
@@
We visualize the embeddings of $M=300$ temporal tokens to investigate their distribution in embedding space. We employ PCA (Abdi and Williams, 2010) to reduce the dimensionality of the temporal tokens to 1D, 2D, and 3D representations. For all reductions, we use the reduced values as coordinates, incorporating a gradient color scheme in which the color of the data points changes progressively with the token index, as illustrated in Figure 5. Our observations reveal that temporal tokens with similar indices tend to cluster together, exhibiting a continuous transition from tokens with smaller indices (light colors) to those with larger indices (darker colors).
A.7 More experiment results
As shown in Table 10, we incorporate a subset of the training sets of Charades-STA and ActivityNet-Captions into the 3rd training stage and re-train the model from the checkpoint of the 2nd training stage, which achieves better performance. This also greatly improves the performance on NExT-GQA as illustrated in Table 11.
@@ A.7 More experiment results
System: You are a good question generator. I need your help in generating question-answer pairs pertaining to the visual event descriptions. I have a video and I will provide you with descriptions of certain segments and their corresponding timestamps within this video. You need to consider these segments comprehensively based on the given description and timestamps and select one segment which you think can provide a HIGH-QUALITY QUESTION. Based on the description of that segment, ask a question related to that segment, as well as one correct answer. Both the proposed answer and question should be consistent with the content of the give description. BE CAREFUL! Your proposed questions and answers should follow these rules: (0) Avoid choosing the segment spanning across the whole video. (1) The question you raised should include causal and temporal relationships as much as possible. Question types should be diverse including WHY, HOW, WHAT, WHERE, etc. (2) NEVER involve anything that is not covered in the given descriptions. (3) The answer should NEVER appear in your question. (4) Your answer should be a phrase no more than 7 words. Keep your answers concise and accurate. Demonstrations: User: video duration: 82.73 seconds segment-1: [0.83, 19.86] A young woman is seen standing in a room and leads into her dancing. segment-2: [17.37, 60.81] The girl dances around the room while the camera captures her movements. segment-3: [56.26, 79.42] She continues dancing around the room and ends by laying on the floor. Response: chosen segment: segment-3 segment timestamps: [56.26, 79.42] question: What did the girl do after she ended dancing? answer: lay on the floor $\cdots$ (other in context demonstrations) $\cdots$
@@
Prompts for Video-Caption Alignment: (1) "Describe the following video concisely.", (2) "Provide a brief description of the given video clip.", (3) "Offer a succinct explanation of the footage presented.", (4) "Summarize the visual content of the following video.", (5) "Give a short and clear explanation of the subsequent video clip.", (6) "Share a concise interpretation of the video provided.", (7) "Present a compact description of the clip’s key features.", (8) "Relay a brief, clear account of the video shown.", (9) "Render a clear and concise summary of the video below.", (10) "Write a terse but informative summary of the following video clip." Prompts for Temporal Sentence Grounding: (1) "When does <query><query> happen in the video?", (2) "At what time does the occurrence <query><query> take place in the video?", (3) "During which part of the video does <query><query> occur?", (4) "When in the video does the <query><query> incident occur?", (5) "At which moment does <query><query> take place in the video?", (6) "During which phase of the video does <query><query> happen?", (7) "When does the <query><query> event occur in the video?", (8) "At what time does <query><query> occur in the video sequence?", (9) "When does the <query><query> situation take place in the video?", (10) "At which time interval in the video can we see <query><query> occurring?" Prompts for Dense Video Captioning: (1) "Localize a series of activity events in the video, output the start and end timestamp for each event, and describe each event with sentences.", (2) "Detect and report the start and end timestamps of activity events in the video, along with descriptions.", (3) "Pinpoint the time intervals of activity events in the video, and provide descriptions for each event.", (4) "Can you compile a list of the activities and their timestamps featured in the video?", (5) "I need you to scrutinize the video and catalog every event it contains, along with the timestamps." Prompts for Temporal Referring: (1) "What is happening from <start><start> to <end><end>?", (2) "What is taking place between <start><start> and <end><end>?", (3) "What events unfold between <start><start> and <end><end>?", (4) "What is happening during the period from <start><start> to <end><end>?", (5) "What occurs between <start><start> and <end><end>?", (6) "What is going on from <start><start> to <end><end>?", (7) "How do things progress from <start><start> to <end><end>?", (8) "Can you describe what happens from <start><start> to <end><end>?", (9) "Describe the events occurring between <start><start> and <end><end>.", (10) "Narrate the actions that unfold from <start><start> to <end><end>."